OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9 8. [6.03] What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9

OpenStudy (tkhunny):

Have you considered showing your work? For the first, I might be tempted to multiply the second equation by 2 and then add the two equations.

OpenStudy (lalipapa123):

1.Substitue the value of x to the 2nd eq'n Substitute the value of y to the 1st eq'n the value you get will be 9. 2. Standard form: 4x+3y=10; 4x+5y=14; substitute/eliminate x = -3/4y+5/2 substitute/eliminate y = +2 Solution: x = 1; y = 2; (x, y) = (1, 2) Standard form: 5x+4y=1; 4x+3y=-1; substitute/eliminate x = -4/5y+1/5 substitute/eliminate y = +9 Solution: x = -7; y = 9; (x, y) = (-7, 9) hope this may help you if so plzzz medal me!